# Progressive Autoregressive Video Diffusion Models

### v1

@@ 3.2 Autoregressive Generation
Variable Length
@@
The above design only allows for autoregressive video extension given an initial video of length $F$. In addition, the noisy latents remaining in the attention window ${\mathbf{x}}_{0:F-1}^{\bm{\tau}^{\prime}_{1:S}}$ (line 8 of Alg. 1) are discarded after the end of the autoregressive inference, which can cause wasted computing resources and inaccurate handling of the ending of text prompt. To enable text-to-long-video generation without starting latents and proper generation of ending latents, we extend the base design in eqs. 8 and 1 to add an initialization stage and an termination stage, where the model operates on variable attention window lengths from $1$ to $F-1$. During initialization, we simply disable the “removing ${\textnormal{x}}_{0}^{0}$” operation in line 8 of Alg. 1: starting from a noisy latent ${\textnormal{x}}_{0}^{T}$, we denoise it for one step and append a new noisy latent to obtain $\{{\textnormal{x}}_{0}^{\tau^{\prime}_{S-1}},{\textnormal{x}}_{1}^{T}\}$; we repeat this by $F-1$ times to obtain ${\mathbf{x}}^{\bm{\tau}^{\prime}_{1:S}}_{0:F-1}=\left\{{\textnormal{x}}_{1}^{\tau^{\prime}_{1}},\ldots,{\textnormal{x}}_{F-2}^{\tau^{\prime}_{S-1}},{\textnormal{x}}_{F-1}^{T}\right\}$, i.e. the input to line 5 of Alg. 1. During termination, we disable the “append ${\textnormal{x}}_{F-1}^{T}$” operation: starting with $F$ latents ${\mathbf{x}}^{\bm{\tau}^{\prime}_{1:S}}_{0:F-1}=\left\{{\textnormal{x}}_{0}^{\tau^{\prime}_{1}},\ldots,{\textnormal{x}}_{F-2}^{\tau^{\prime}_{S-1}},{\textnormal{x}}_{F-1}^{T}\right\}$, we denoise them for one step, save and remove the first latent ${\textnormal{x}}_{0}^{0}$ and shift the latents by $-1$ to obtain ${\mathbf{x}}^{\bm{\tau}^{\prime}_{1:S-1}}_{0:F-2}=\left\{{\textnormal{x}}_{0}^{\tau^{\prime}_{1}},\ldots,{\textnormal{x}}_{F-2}^{\tau^{\prime}_{S-1}}\right\}$; we repeat this by $F$ times to save and remove all the remaining latents in the attention window. We train the model accordingly on video latent frames with variable lengths ranging from $1$ to $F-1$, following the noise levels described above.
@@ 3.3 Training
While regular video diffusion models always treat the latent frames as a single entity by assigning a single noise level t to all of them, we extend this formulation by modeling each frame independently and assigning them with varying noise levels ${\mathbf{t}}_{0:F-1}$. This change in the noise level distribution typically requires finetuning the pre-trained video diffusion model to adapt to our progressive noise level distribution. We finetune pre-trained video diffusion models by modifying the noise levels during training. Regular diffusion model training (Ho et al., 2020; Liu et al., 2022; 2023) involves uniformly sampling a noise level ${\textnormal{t}}\in[0,T)$, adding noise to the samples ${\mathbf{x}}^{0}_{0:F-1}$ via the forward diffusion process (eqs. 1 and 2), and computing the loss (eq. 3). To enable sampling with progressive noise levels in eqs. 7 and 8, we simply switch to per-frame training noise levels ${\mathbf{t}}_{0:F-1}$. In our experiment, we observed that, similar to the sampling noise levels $\bm{\tau}^{\prime}_{0:S}$ in eq. 7, training on a simple linear noise schedule yielded satisfactory results for all reported experiments. During training, the noise levels ${\bm{t}}$ is perturbated by a random shift $\delta$ to preserve the coverage of the full diffusion timestep range $[0,T)$ (Song et al., 2020a). $\delta=0.4\epsilon({\textnormal{t}}_{i}-{\textnormal{t}}_{i+1}),\bm{\epsilon}\sim\mathcal{N}(0,{\bm{I}})$ is randomly sampled for each training iteration and remains constant for all ${\mathbf{t}}_{0:F-1}$ within that iteration.
@@ 4 Experiments
In this section, we follow our terminology as discussed and defined in Sec. 2.2, including the replacement-with-noise method, the replacement-without-noise method, and attention window.
@@ 4.1 Implementation
Our models and baseline models
We implement our progressive autoregressive video diffusion models by fine-tuning from pre-trained models. Specifically, we use two video diffusion models based on the diffusion transformer architecture (Peebles & Xie, 2023; Brooks et al., 2024): Open-Sora v1.2 (Zheng et al., 2024) (denoted as O) and a modified variant of Open-Sora (denoted as M in later experiments). Both models are latent video diffusion models (Blattmann et al., 2023b), each utilizing a corresponding 3D VAE that encodes 17 (O) or 16 (M) video frames into 5 latent representations. O generates videos at 240$\times$424 resolution 24 FPS with 30 sampling steps. M produces results at 176$\times$320 resolution 24 FPS with 50 sampling steps. Based on O and M, we also implement two baseline autoregressive video generation methods, replacement-with-noise (denoted as RW) and replacement-without-noise (denoted as RN) (Sec. 2.2), to compare with our proposed progressive autoregressive (denoted as PA) video generation method (Sec. 3).
@@
O undergoes masked pre-training (Zheng et al., 2024), where the masked latents ${\mathbf{x}}_{0:E-1}^{0}$ are clean without any added noise (Zheng et al., 2024). This allows the O base model to perform autoregressive video generation with the replacement-without-noise method. We denote this model as RN-O-base. Such training also allows O to learn that the noise levels ${\mathbf{t}}_{0:F-1}$ can be independent with respect to the latent frames and thus enables our progressive autoregressive video denoising sampling procedure (Secs. 3.2 and 1) to work training-free. We denoise this model as PA-O-base.
Training details
We train M and O on captioned image and video datasets, containing 1 million videos and 2.3 billion images. These data are licensed and have been filtered to remove low-quality content. We train PA-M on video clips of $16,32,...,176$ frames that correspond to $F=5,10,...,55$ latents. The $F=55$ attention window length is derived by setting $F=S+5$, where $S=50$ is the number of sampling steps in M ($S=30$ in O) and $5$ is the length of an additional chunk of latent frames. The advantage of such design is described in the below “Keeping clean frames available in temporal self-attention” paragraph. The shorter latent frame lengths $F=5,10,...,50$ are used for the variable length training, as discussed in Sec. 3.2. RW-M is trained on videos of $64$ frames that corresponds to $F=20$ latents.
Modification to the base model
To implement progressive autoregressive video diffusion models on top of their pre-trained foundation video diffusion models, we do not need to modify the base model architectures. Instead, we only need to modify the model’s forward, training, and inference procedures. In the training and inference procedures, we replace the single noise level ${\textnormal{t}}\in[0,T)$ from regular diffusion model training (Ho et al., 2020; 2022b) with our per-frame noise level ${\mathbf{t}}_{0:F-1}$ and $\bm{\tau}^{\prime}_{0:S-1}$ (Secs. 3.3 and 3.1). To accommodate this change, we only need to make a single modification to the the noise level embedding computation in the model’s forward procedure. While the regular timestep only has the batch size dimension $B$, our progressive timesteps has two dimensions $B,F$. We first flatten them into the batch dimension of size $B\times F$, pass it to the timesteps embedding module, unflatten the two dimensions, and finally broadcast the timestep embedding to the same shape of the latents so they can be combined through either addition, concatenation, modulation, or cross-attention (Perez et al., 2018; Vaswani et al., 2017; Peebles & Xie, 2023).
Chunk-by-chunk denoising
3D VAE (Zheng et al., 2024) usually encode and decode video latent frames chunk-by-chunk. In our early experiments on PA-M and PA-O with $F=S$, we find that there is serious cumulative error when all latents are given different noise levels and the attention window is shifted by one latent at a time, causing the generated videos to diverge quickly after a few seconds, as shown in Ablation 2 in Fig. 5. We resolve the problem by treating each chunk of latents as a single latent: they are assigned with the same noise level, and are added and removed from the attention window together. In other words, for the 3D VAE chunk size of $C=5$ latents in M and O, we shift the attention window by $C$ latents every $C$ sampling steps. Effectively, the $C$ latents that belong to the same chunk always have the same noise level $t$ and are added to or removed from the attention window together. Our ablation experiments on both O and M show that the chunked training and inference substantially improves the generation result.
Keeping clean frames available in temporal self-attention
The default design of our progressive autoregressive video denoising process presented in Sec. 3.2 results in temporal jittering. This is because the clean latents ${\mathbf{x}}_{0:C-1}^{0}$ are immediately removed from the attention window; as the later latents cannot attend to the previous clean latents, it is hard for the model to denoise the later latents to be perfectly temporally consistent with the previous clean latents. In practice, we always keep a chunk of $C$ clean latents, e.g., $C=5$ mentioned ahead in our training details, in front of all the noisy latents, extending the attention window lengths $F$ from $50$ to $55$ and from $30$ to $35$ for M and O respectively. This helps resolving the frame-to-frame discontinuity issue.
@@ 4.2 Long video generation
Baselines
As discussed in Sec. 4.1, using our base models, we implement two baseline autoregressive video generation methods on three models, which are denoted as RW-M, RN-O-base, and RN-O. We also compare to Stable Video Diffusion (SVD) (Blattmann et al., 2023a) and StreamingT2V (Henschel et al., 2024) model families. Specifically, we consider the SVD-XT model from SVD, a image-to-video model that generates a short video clip of 25 frames at 576x1024 resolution given an conditioning image. We apply it autoregressively, using the last image of the previous clip as the condition for generating a new clip. This is equivalent to the replacement-without-noise method except that it only conditions on a single frame rather than a chunk of 17 frames as RN-O. We also consider the StreamingSVD model from StreamingT2V, a image-to-long-video generation model that uses SVD as the base model (Henschel et al., 2024); its autoregressive video generation is enabled by training additional modules that connect to the base model via cross-attention. Similar to our progressive autoregressive video diffusion models, StreamingSVD can autoregressively generate long videos at 720x1280 resolution with arbitrary lengths, which we set to 1440 frames.
Testing set
Our testing set consists of 40 text prompts and the corresponding real videos, sampled from Sora (Zheng et al., 2024) demo videos, MiraData (Ju et al., 2024), UCF-101 (Soomro, 2012), and LOVEU (Wu et al., 2023a; b). For each text prompt, we generate two videos with 1440 frames, 60 seconds long at 24 FPS, resulting in a total of 80 videos. We use these 80 videos from each model for both quantitative and qualitative results, unless specified otherwise. Due to computation resource limitations of sampling 1-minute long videos, we only obtained partial results from M-PA and StreamingSVD, including 48 videos from 24 text prompts and 40 videos from 40 text prompts respectively. This testing set measures the zero-shot long video generation ability of the models, since none of them are specifically trained on any of the above datasets.
Since our focus is on long video generation, we focus on the video extension capability of the models rather than the text-to-short-video generation capability. Thus, we use the initial frames of the videos as the condition for all models, similar to the setting in (Henschel et al., 2024). M, O (Zheng et al., 2024), StreamingSVD (Henschel et al., 2024), and SVD-XT (Blattmann et al., 2023a) use 16, 17, 1, and 1 frames from the real video as the initial condition.
Metrics
@@
We consider 6 metrics in VBench (Huang et al., 2024): subject consistency, background consistency, motion smoothness, dynamic degree, aesthetic quality, and imaging quality. We compute average metrics using VBench-long, where each metric is computed on 30 2-second clips for each 60-second video; for subject and background consistency, a clip-to-clip metric is considered in addition to the average metric over the clips. We also show how the metrics vary over time by plotting the metrics over the 30 2-second clips averaged over the 80 60-second videos. Similar to Henschel et al. (2024), we also use the Adaptive Detector algorithm from PySceneDetect (PySceneDetect, ) to count the number of detected scene changes, where Num Scenes $=1$ means that there is no scene change detected.
Quantitative Results
We present the average metrics for each model in Tab. 1. The metrics are averaged over all the videos that each model generates from our testing set described above. In Fig. 3, we also illustrate the trend of metrics over the 1-minute duration of videos for each model. We find that subject consistency, background consistency, and motion smoothness metrics are not discriminative, as all models have obtained similar results on these metrics on average and can also maintain them over the 1-minute duration. We summarize the quantitative results on the remaining metrics of each model below.
Table 1: Quantitative comparison of our progressive autoregressive video generation (PA) and two baseline methods replacement-with-noise (RW) and replacement-without-noise (RN) on two base models (M and O), StreamingT2V (Henschel et al., 2024), and Stable Video Diffusion (SVD) (Blattmann et al., 2023a).
Subject Background Motion Dynamic Aesthetic Imaging Num
Consistency $\uparrow$ Consistency $\uparrow$ Smoothness $\uparrow$ Degree $\uparrow$ Quality $\uparrow$ Quality $\uparrow$ Scenes $\downarrow$
PA-M (ours) 0.7923 0.8964 0.9896 0.8000 0.4726 0.5927 1.75
RW-M 0.8001 0.8851 0.9836 0.3958 0.4123 0.5961 1.10
PA-O-base (ours) 0.7656 0.8880 0.9859 0.5625 0.4582 0.5033 2.04
RN-O-base 0.7406 0.8820 0.9873 0.5750 0.4034 0.4464 5.19
StreamingSVD 0.8172 0.8916 0.9929 0.65 0.4264 0.5566 1.08
SVD-XT 0.6102 0.8136 0.9724 0.9875 0.3019 0.4814 2.10
Figure 3: VBench (Huang et al., 2024) scores of generated videos over the 60-second duration, averaged over 80 videos from our testing set. The scores are computed on 30 2-second clips. Our models M-PA and O-PA can best maintain the level of dynamic degree, aesthetic quality, and imaging quality over time compared to other baselines. Notably, baselines that use the same model as ours, M-RW and O-RN, both exhibit substantial drop in dynamic degree, aesthetic quality, and imaging quality.
Our PA-M obtains the best overall quality; on average, it has substantially better dynamic degree than all except SVD-XT, the best aesthetic and imaging quality, and only slightly worse number of scenes than RW-M and StreamingSVD; it is able to best maintain the level of dynamic degree, aesthetic and imaging quality over the 1-minute duration.
Although the baseline RW-M has a nearly perfect number of scenes $\approx 1$, it has a poor dynamic degree that degrades substantially after a few seconds, where most of its generated videos are static; furthermore, its aesthetic and image qualities also degrades over the 1-minute duration.
Our PA-O-base achieves slightly better dynamic degree, aesthetic quality and imaging quality on average and better maintains them over time compared to RN-O-base; it also substantially reduces the number of scenes and is effectively generating longer video shots; these results are impressive given that they are obtained training-free.
The other baseline RN-O-base have worse aesthetic and imaging quality than our PA-O-base both on average and over time; it also has the worst number of scenes, making it fail at the core goal of long video generation.
StreamingSVD achieves good overall quality; it achieves the best number of scenes, decent dynamic degree, aesthetic quality and imaging quality and can maintain them over time; however, they worse than our PA-M on average.
SVD-XT, although achieves the best dynamic degree, also has the worst aesthetic and imaging quality both on average and over time, making its results undesirable.
Qualitative Results

### v2

@@ 3 Progressive Autoregressive Video Diffusion Models
We consider long video generation with video diffusion models. As discussed in Sec. 2.2, existing video diffusion models can only generate short video clips up to a limited length $F$, and the replacement methods [15, 58, 7] suffer from various flaws. We describe a more natural formulation of autoregressive long video generation, which we call Progressive Autoregressive Video Diffusion Models (PA-VDM). We propose a per-frame progressively increasing noise schedule, which is inspired by [4]. During training, we finetune pre-trained video diffusion models to adapt to our noise schedule; during sampling, our models adopt such noise schedule and autoregressively generate video frames.
3.1 Progressive Noise Levels and Autoregressive Generation
Conventional video diffusion methods assign a single noise level $t$ to all the latent frames. Inspired by [4], we adopt per-frame noise levels ${\mathbf{t}}_{0:F-1}=\{t_{0},t_{1},...,t_{F-1}\}$ to the $F$ latent frames in the attention window. In particular, we consider monotonically increasing noise levels for each frame, where earlier frames are less noisy and later frames are more noisy. In this work, we consider the linear sampling noise schedule with $S$ sampling steps
$\displaystyle\bm{\tau}_{0:S}=\left\{0,\frac{T}{S},\frac{2T}{S},\ldots,\frac{(S-1)T}{S},T\right\}$ (7)
which is monotonically increasing. Given a sampling noise schedule, instead of all the frames sharing a noise level and jointly going through the schedule as in conventional video diffusion models, each frame now goes through the schedule independently; at each step, the per-frame noise levels $\bm{\tau}$ still maintain the progressively increasing pattern.
Since both the sampling noise schedule and our target per-frame noise levels are monotonically increasing, we can now set per-frame noise levels ${\mathbf{t}}_{0:F-1}$ to be an interpolation of the sampling noise schedule $\bm{\tau}$. Let us first consider the simple case of $F=S$, when our per-frame progressive noise levels can equal to either ${\mathbf{t}}=\bm{\tau}_{0:S-1}$ or ${\mathbf{t}}=\bm{\tau}_{1:S}$. At each sampling step, the video diffusion model takes $\bm{\tau}_{0:S-1}$ as input and predicts $\bm{\tau}_{1:S}$
$\displaystyle p_{\theta}\left({\textnormal{x}}_{0}^{\tau_{0}},{\textnormal{x}}_{1}^{\tau_{1}},...,{\textnormal{x}}_{F-2}^{\tau_{S-2}},{\textnormal{x}}_{F-1}^{\tau_{S-1}}\middle|{\textnormal{x}}_{0}^{\tau_{1}},{\textnormal{x}}_{1}^{\tau_{2}},...,{\textnormal{x}}_{F-2}^{\tau_{S-1}},{\textnormal{x}}_{F-1}^{\tau_{S}}\right)$ (8)
We illustrate progressive noise levels when $F=S$ in Fig. 2.
Now we construct our autoregressive generation algorithm for video latent frames with progressive noise levels. Notice that the input and output noise levels in Eq. 8, $\bm{\tau}_{0:S-1}$ and $\bm{\tau}_{1:S}$, only differ by $\tau_{0}=0$ and $\tau_{S}=T$. We can simply transition the output frames back into the correct input noise levels by removing the clean frame ${\textnormal{x}}_{0}^{0}$ at the front, shifting the frame sequence forward by one frame, and appending a new noisy frame ${\textnormal{x}}_{F-1}^{T}\sim\mathcal{N}(\mathbf{0},{\bm{I}})$ at back, as illustrated in Fig. 1. We describe the autoregressive generation algorithm when $F=S$ in Alg. 1. The algorithm requires a clean short video $\bm{x}_{0:F-1}^{0}$ as initialization and extends from it. We describe how to avoid this requirement in Sec. 3.2.
More generally, when $F$ is a multiple of $S$, e.g. $F=90,S=30$, every set of $F/S=3$ frames would always share the same noise level during denoising and be removed from the attention window together as they reach $t=0$; when $S$ is a multiple of $F$, e.g. $F=10,S=30$, the save, shift, and append operations for the sequence of frames (line 6, 7, 8 in Alg. 1) would happen once every $S/F=3$ steps.
Note that, regardless of what the noise level a frame initially has, it always goes through the same noise level schedule $\bm{\tau}_{S:0}$ as in conventional diffusion models. Thus, for each individual frame, it is still modeled under the valid assumptions in diffusion model training [13, 25, 26, 27] and sampling [42]. We only diverge from the noise level assumption in conventional video diffusion models [15]: now, each frame is modeled independently instead of jointly with the whole sequence of frames, and the progressive autoregressive video diffusion model attends to frames with different noise levels ${\mathbf{t}}_{0:F-1}$ instead of the same noise level $t$. Thus, we can obtain our progressive autoregressive video diffusion models from pre-trained video diffusion models by adapting the model to the new noise level distribution through finetuning. This saves us from the highly demanding computation cost of video diffusion model pre-training [34, 55].
Intuitively, the benefit of our progressive video denoising process is that it gradually establishes correlation among consecutive latent frames. Given some existing video frames as conditioning, it is challenging for video diffusion models to produce temporally consistent extension frames from newly sampled noisy frames [36]. In contrast to the replacement-with-noise method [1, 15] where the frames are denoised together at the same noise level, our progressive video denoising encourages the later frames with higher uncertainty to follow the patterns of the earlier and more certain frames, facilitating modeling a smoother temporal transition and better preserving motion velocity. Compared to the replacement-without-noise method where there is a large noise level gap between the clean condition frames $\bar{{\mathbf{x}}}_{0:E-1}^{0}$ and the noisy frames ${\mathbf{x}}_{E:F-1}^{\tau_{i}}$, our method provides smoother attention correspondence, where the difference between neighboring noise levels is only $\frac{T}{S}$, as illustrated in Eqs. 7 and 2.
3.2 Variable Length
@@
The above design only allows for autoregressive video extension given an initial video of length $F$. In addition, the noisy frames remaining in the attention window ${\mathbf{x}}_{0:F-1}^{\bm{\tau}_{1:S}}$ (line 8 of Alg. 1) are discarded after the end of the autoregressive inference, which can cause wasted computing resources and inaccurate handling of the ending of text prompt. To enable text-to-long-video generation without any starting condition frames and properly ending the generation without wasting computation, we extend the base design in Eqs. 8 and 1 to add an initialization stage and a termination stage, where the model operates on variable attention window lengths from $1$ to $F-1$. During initialization, we simply disable the “removing ${\textnormal{x}}_{0}^{0}$” operation in line 8 of Alg. 1: starting from a noisy frame $\{{\textnormal{x}}_{0}^{T}\}$, we denoise and append to obtain $\{{\textnormal{x}}_{0}^{\tau_{S-1}},{\textnormal{x}}_{1}^{T}\}$; we repeat this by $F-1$ times to obtain ${\mathbf{x}}^{\bm{\tau}_{1:S}}_{0:F-1}=\left\{{\textnormal{x}}_{1}^{\tau_{1}},\ldots,{\textnormal{x}}_{F-2}^{\tau_{S-1}},{\textnormal{x}}_{F-1}^{T}\right\}$, i.e. the input to line 5 of Alg. 1. During termination, we disable the “append ${\textnormal{x}}_{F-1}^{T}$” operation in line 6 and 7 of Alg. 1: starting with $F$ frames ${\mathbf{x}}^{\bm{\tau}_{1:S}}_{0:F-1}=\left\{{\textnormal{x}}_{0}^{\tau_{1}},\ldots,{\textnormal{x}}_{F-2}^{\tau_{S-1}},{\textnormal{x}}_{F-1}^{T}\right\}$, we denoise, save and remove to obtain ${\mathbf{x}}^{\bm{\tau}_{1:S-1}}_{0:F-2}=\left\{{\textnormal{x}}_{0}^{\tau_{1}},\ldots,{\textnormal{x}}_{F-2}^{\tau_{S-1}}\right\}$; we repeat this by $F$ times to save and remove all the remaining frames in the attention window. We train the model accordingly on video latent frames with variable lengths ranging from $1$ to $F-1$, following the noise levels described above.
3.3 Chunked Frames
3D VAEs [20, 57, 34, 58] usually encode and decode video latent frames chunk-by-chunk. In our early experiments, we find that naively implementing our method on latent video diffusion models, i.e. when all latent frames are given different noise levels and the attention window is shifted by one frame at a time, leads to serious cumulative error and the videos diverge quickly after a few seconds, as shown in Ablation 2 in Fig. 6. We resolve the problem by treating a chunk of latent frames as a whole: they are assigned with the same noise level, and are added and removed from the attention window together. In other words, for a 3D VAE chunk size of $C$ latent frames, e.g. $C=5$ as mentioned in Sec. 4, we shift the attention window by $C$ frames every $C$ sampling steps. Effectively, the $C$ frames that belong to the same chunk always have the same noise level $t$ and are added to or removed from the attention window together. Our ablation experiments shows that, for models using a 3D VAE, treating a chunk of frames as a whole effectively prevents accumulated errors that would lead to divergence.
3.4 Overlapped Conditioning
In our early experiments, naively implementing our method on video diffusion models results in temporal jittering. We hypothesize that this is because the clean frames ${\mathbf{x}}_{0:C-1}^{0}$ are immediately removed from the attention window; as the later frames cannot attend to the previous clean frames, it is hard for the model to denoise the later frames to be perfectly temporally consistent with the previous clean frames. In practice, we always keep a chunk of $C$ clean frames by prepending it to the attention window. Our ablation study shows that overlapped conditioning helps resolving the frame-to-frame discontinuity issue.
Overlapped conditioning requires an additional inference cost at $C/F$ ($5/50$ in our implementation) of the original cost. When using the same number of conditioning frames $E$ and $F$, the replacement methods [15, 58, 7] and ours have the same inference efficiency. The key advantage of our method is that the large overlap of noisy frames enables the model to preserve the high-level information—such as motion—from prior frames. Thus, we only need a single chunk of $C$ clean condition frames to propagate high-frequency details and prevent per-chunk temporal jittering. In contrast, the replacement methods need to balance the tradeoff between more overlap between video clips or better inference efficiency. In practice, their implementation [58] often use one chunk of frames as condition to save inference computation, but the limited overlap causes unnatural motion transition and abrupt scene changes across clips, as discussed in Sec. 4.2.
@@ 3.5 Training
As described in Sec. 3.1, PA-VDM requires change in the noise level distribution. We finetune pre-trained video diffusion model to adapt to our progressive noise level distribution. Conventional diffusion model training [13, 26, 27] involves uniformly sampling a noise level ${\textnormal{t}}\in[0,T)$, adding noise to the samples ${\mathbf{x}}^{0}_{0:F-1}$ via the forward diffusion process (Eqs. 1 and 2), and computing the loss (Eq. 3). During the finetuning process for PA-VDM, we simply continue with the conventional video diffusion model training but with our per-frame progressive training noise levels ${\mathbf{t}}_{0:F-1}$. In our experiment, we observed that, similar to the sampling noise levels $\bm{\tau}_{0:S}$ in Eq. 7, training on a simple linear noise schedule yielded satisfactory results for all reported experiments. During training, the noise levels ${\mathbf{t}}$ is perturbated by a random shift $\delta$ to fully cover of the diffusion timestep range $[0,T)$ [41]. $\delta=0.4\epsilon({\textnormal{t}}_{i}-{\textnormal{t}}_{i+1}),\bm{\epsilon}\sim\mathcal{N}(0,{\bm{I}})$ is randomly sampled for each training iteration and remains constant for all ${\mathbf{t}}_{0:F-1}$ within that iteration.
@@ 4 Experiments
Table 1: Quantitative comparison of our progressive autoregressive video generation (PA) and two baseline methods replacement-with-noise (RW) and replacement-without-noise (RN) on two base models (M and O), and other baselines StreamingT2V [12], Stable Video Diffusion (SVD) [1], and FIFO-Diffusion [19].
Subject Background Motion Dynamic Aesthetic Imaging Num
Consistency $\uparrow$ Consistency $\uparrow$ Smoothness $\uparrow$ Degree $\uparrow$ Quality $\uparrow$ Quality $\uparrow$ Scenes $\downarrow$ FVD $\downarrow$
PA-M (ours) 0.7923 0.8964 0.9896 0.8000 0.4726 0.5927 1.75 358.020
RW-M 0.8001 0.8851 0.9836 0.3958 0.4123 0.5961 1.10 669.747
PA-O-base (ours) 0.7656 0.8880 0.9859 0.5625 0.4582 0.5033 2.04 548.117
RN-O-base 0.7406 0.8820 0.9873 0.5750 0.4034 0.4464 5.19 600.690
StreamingSVD 0.8172 0.8916 0.9929 0.65 0.4264 0.5566 1.08 440.272
SVD-XT 0.6102 0.8136 0.9724 0.9875 0.3019 0.4814 2.10 702.343
FIFO-OSP 0.7577 0.8990 0.9731 0.75 N/A 0.5675 18.32 975.459
@@ 4.1 Implementation
Our models and baseline models We implement our progressive autoregressive video diffusion models by fine-tuning from pre-trained models. Specifically, we use two video diffusion models based on the diffusion transformer architecture [32, 3]: Open-Sora v1.2 [58] (denoted as O) and a modified variant of Open-Sora (denoted as M in later experiments). Both models are latent video diffusion models [2], each utilizing a corresponding 3D VAE that encodes 17 (O) or 16 (M) raw video frames into 5 latent frames. O generates videos at 240$\times$424 resolution 24 FPS with 30 sampling steps. M produces results at 176$\times$320 resolution 24 FPS with 50 sampling steps. Based on O and M, we also implement two baseline autoregressive video generation methods, replacement-with-noise (denoted as RW) and replacement-without-noise (denoted as RN) (Sec. 2.2), to compare with our proposed progressive autoregressive (denoted as PA) video generation method (Sec. 3).
@@
O undergoes masked pre-training [58], where the masked latent frames ${\mathbf{x}}_{0:E-1}^{0}$ are clean without any added noise [58]. This allows the O base model to perform autoregressive video generation with the replacement-without-noise method. We denote this model as RN-O-base. Such training also allows O to learn that the noise levels ${\mathbf{t}}_{0:F-1}$ can be independent with respect to the latent frames and thus enables our progressive autoregressive video denoising sampling procedure (Alg. 1) to work training-free. We denote this model as PA-O-base. Please refer to Appendix E for training details.
@@ 4.2 Long video generation
The baseline methods are described in Sec. F.1.
@@
Metrics We consider 6 metrics in VBench [17]: subject consistency, background consistency, motion smoothness, dynamic degree, aesthetic quality, and imaging quality. We compute average metrics using VBench-long, where each metric is computed on 30 2-second clips for each 60-second video; for subject and background consistency, a clip-to-clip metric is considered in addition to the average metric over the clips. We also show how the metrics vary over time by plotting the metrics over the 30 2-second clips averaged over the 80 60-second videos.
Similar to [12], we also use the Adaptive Detector algorithm from PySceneDetect [35] to count the number of detected scene changes, where Num Scenes $=1$ means that there is no scene change detected.
We also compute Fréchet Video Distance (FVD) [47] to measure the overall quality of the generated videos compared to real videos. We adopt the improved implementation of FVD proposed in [8] using the VideoMAE-v2 [50] model. The FVD metric usually requires a large number of video samples in order to produce a reliable value. Since our testing set includes only 40 real videos and each model only generate 80 videos, naively computing FVD on them results in erroneous values such as -3.62e+64. Instead, we compute FVD on the 2-second clips of the long videos, so that we have 1495 real videos and 2400 generated videos.
Quantitative Results We present the average metrics for each model in Tab. 1. The metrics are averaged over all the videos that each model generates from our testing set described above. Our PA-M has the best results overall. Notably, it surpasses other methods in FVD by a substantial margin, illustrating that its results are the most realistic. It also achieves either the best or close-to-best in other metrics. Its replacement-with-noise counterpart, RW-M, suffers from poor Dynamic Degree and FVD, because its videos are mostly static. Our RW-O-base surpasses its replacement-without-noise counterpart RN-O-base in all metrics except for being close at Dynamic Degree, while using the exact same model parameters without any finetuning. RN-O-base mainly suffers from a high number of scene changes.
In Fig. 3, we also illustrate the trend of metrics over the 1-minute duration of videos for each model. Our models M-PA and O-PA can best maintain the level of all metrics, while their replacement-method counterparts, M-RW and O-RN, both exhibit distinct reduction in dynamic degree, aesthetic quality, and imaging quality.
Figure 3: VBench [17] scores over the 60-second duration, which are computed on 30 2-second clips.
Figure 4: Human evaluation results comparing long video methods on long-shot (L), motion (M), temporal consistency (C), and overall (O).